# Test-time Computing: from System-1 Thinking to System-2 Thinking

### v1

@@ Abstract
The remarkable performance of the o1 model in complex reasoning demonstrates that test-time computing scaling can further unlock the model’s potential, enabling powerful System-2 thinking. However, there is still a lack of comprehensive surveys for test-time computing scaling. We trace the concept of test-time computing back to System-1 models. In System-1 models, test-time computing addresses distribution shifts and improves robustness and generalization through parameter updating, input modification, representation editing, and output calibration. In System-2 models, it enhances the model’s reasoning ability to solve complex problems through repeated sampling, self-correction, and tree search. We organize this survey according to the trend of System-1 to System-2 thinking, highlighting the key role of test-time computing in the transition from System-1 models to weak System-2 models, and then to strong System-2 models. We also point out a few possible future directions. [^1] https://github.com/Dereck0602/Awesome_Test_Time_LLMs.
Test-time Computing: from System-1 Thinking to System-2 Thinking
Yixin Ji1, Juntao Li1thanks:   Corresponding author., Hai Ye2, Kaixin Wu3, Jia Xu3, Linjian Mo3, Min Zhang1 1School of Computer Science and Technology, Soochow University 2Department of Computer Science, National University of Singapore 3Ant Group jiyixin169@gmail.com; {ljt,minzhang}@suda.edu.cn yehai@comp.nus.edu.sg; {daniel.wkx,steve.xuj,linyi01}@antgroup.com
@@ 1 Introduction
Over the past decades, deep learning with its scaling effects has been the driving engine behind the AI revolution. Particularly in the text modality, large language models (LLMs) represented by the GPT series (Radford et al., 2018, 2019; Brown et al., 2020; Ouyang et al., 2022; OpenAI, 2023) have demonstrated that larger models and more training data lead to better performance on downstream tasks. However, on the one hand, further scaling in the training phase becomes difficult due to the scarcity of data and computational resources (Villalobos et al., 2024); on the other hand, existing models still perform far below expectations in terms of robustness and handling complex tasks. These shortcomings are attributed to the model’s reliance on fast, intuitive System-1 thinking, rather than slow, deep System-2 thinking (Weston and Sukhbaatar, 2023). Recently, the o1 model (OpenAI, 2024), equipped with System-2 thinking, has gained attention for its outstanding performance in complex reasoning tasks. It demonstrates a test-time computing scaling effect: the greater the computational effort in the inference, the better the model’s performance.
@@
The concept of test-time computing emerged before the rise of LLMs and was initially applied to System-1 models (illustrated in Figure 1). These System-1 models can only perform limited perceptual tasks, relying on patterns learned during training for predictions. As a result, they are constrained by the assumption that training and testing are identically distributed and lack robustness and generalization to distribution shifts (Zhuang et al., 2020). Many works (Wang et al., 2021; Ye et al., 2023) have attempted to improve model robustness with test-time computing, also known as test-time adaptation (TTA). In the era of small models, the mainstream method of TTA is to update model parameters at test-time, so that the model’s representation gradually adapts to the test distribution. For LLMs, due to the high cost of updating parameters, TTA leverages external information to steer model behavior, including modifying inputs, editing representations, and calibrating outputs. By introducing TTA, the System-1 model slows down its thinking process and gradually evolves into a weak System-2 model.
Figure 1: Illustration of test-time computing in the System-1 and System-2 model.
@@
Currently, advanced LLMs with chain-of-thought (CoT) prompting have enabled language models to reach preliminary System-2 thinking, demonstrating the human-like cognitive ability to decompose problems and reason step by step (Wei et al., 2022). However, they still struggle with complex tasks like reasoning and planning (Stechly et al., 2024; Sprague et al., 2024). To achieve stronger System-2 models, researchers employ three types of test-time computing strategies to extend model reasoning’s depth and accuracy: repeated sampling, self-correction, and tree search. Repeated sampling simulates the diversity of human thinking, self-correction enables LLMs to reflect, and tree search enhances reasoning depth.
The remainder of this paper provides a comprehensive survey of the latest research developments of test-time computing in System-1 and System-2 models. In Section 2, we present the background of System-1 and System-2 thinking. Section 3 and Section 4 detail the test-time computing methods for the System-1 and System-2 models. Then, we discuss possible future directions in Section 5.
@@ 2 Background
System-1 and System-2 thinking are psychological concepts (Kahneman, 2011). When recognizing familiar patterns or handling simple problems, humans often respond intuitively. This automatic, fast thinking is called System-1 thinking. In contrast, when dealing with complex problems like mathematical proofs or logical reasoning, deep and deliberate thought is required, referred as System-2 thinking—slow and reflective. In the field of artificial intelligence, researchers also use these terms to describe different types of models (LeCun, 2022). System-1 models respond directly based on internally encoded perceptual information and world knowledge without showing any intermediate decision-making process. In contrast, System-2 models explicitly generate reasoning processes and solve tasks incrementally. Before the rise of LLMs, System-1 models were the mainstream in AI. Although many deep learning models, such as ResNet, Transformer, and BERT, achieve excellent performance in various tasks in computer vision and natural language processing, these System-1 models, similar to human intuition, lack sufficient robustness and are prone to errors. Nowadays, the strong generation and reasoning capabilities of LLMs make it possible to build System-2 models. Wei et al. (2022) propose the CoT, which allows LLMs to generate intermediate reasoning steps progressively during inference. Empirical and theoretical results show that this approach significantly outperforms methods that generate answers directly (Kojima et al., 2022; Zhou et al., 2023; Tang et al., 2024b; Feng et al., 2024a; Li et al., 2024e). However, current System-2 models represented by CoT prompting still have shortcomings. The intermediate processes generated by LLMs may contain errors, leading to cumulative mistakes and ultimately resulting in incorrect answers. Although retrieval-augmented generation (RAG) helps mitigate factual errors (Trivedi et al., 2023; Guan et al., 2024; Wang et al., 2024g; Ji et al., 2024), their impact on improving reasoning abilities remains limited. As a result, CoT-enabled LLMs are still at the weak system-2 thinking stage.
@@
 Auto-CoT (Zhang et al., 2023b), DAWN-ICL (Tang et al., 2024a); etc. , leaf, text width=34em ] ] ] [  Representation
@@
 Calibration [  $k$NN-MT (Khandelwal et al., 2021), AdaNPC Zhang et al. (2023a), Bi-$k$NN (You et al., 2024); etc. , leaf, text width=45.6em ] ] ] [  Test-time
@@
 GenRM (Zhang et al., 2024d), CriticRM (Yu et al., 2024); etc. , leaf, text width=34em ] ] [ Verbal-based [  Liu et al. (2023), LLM-as-a-Judge (Zheng et al., 2023), Auto-J (Li et al., 2024a),
 Prometheus (Kim et al., 2024b, c), Fennec (Liang et al., 2024c); etc. , leaf, text width=34em ] ] ] [  Test-time
@@
 DiVeRSe (Li et al., 2023c), PRS (Ye and Ng, 2024)Zhang et al. (2024e); etc. , leaf, text width=34em ] [  Improvement training: ReST (Gulcehre et al., 2023), vBoN (Amini et al., 2024),
 BoNBoN(Gui et al., 2024), BOND (Sessa et al., 2024)Chow et al. (2024); etc. , leaf, text width=34em ] ] [ Self-correction [  Self-debug (Chen et al., 2024c), RIC (Kim et al., 2023), Critic (Gou et al., 2024),
 Shepherd (Wang et al., 2023b), MAD (Liang et al., 2024b), IoE (Li et al., 2024c),
 Refiner (Paul et al., 2024), Reflexion (Shinn et al., 2023), Du et al. (2023); etc. , leaf, text width=34em ] [  Improvement training: GLoRe (Havrilla et al., 2024), SCoRe(Kumar et al., 2024),
 Self-correct (Welleck et al., 2023), Qu et al. (2024)Zhang et al. (2024g); etc. , leaf, text width=34em ] ] [ Tree Search [  ToT (Yao et al., 2023), RAP (Hao et al., 2023), rStar (Qi et al., 2024),
@@
 MCTS-DPO (Xie et al., 2024), Zhao et al. (2024b)Zhang et al. (2024f); etc. , leaf, text width=34em ] ] ] ] [  Future
 Directions (§5) [  Generalization [  Jia (2024), GRM (Yang et al., 2024), DogeRM (Lin et al., 2024b), Weak-to-strong (Burns et al., 2023); etc. , leaf, text width=45.6em ] ] [  Multi-modal [  MM-CoT (Zhang et al., 2024h), VoT (Wu et al., 2024b), Lee et al. (2024), LLaVA-CoT (Xu et al., 2024); etc. , leaf, text width=45.6em ] ] [  Efficient [  Damani et al. (2024), OSCA (Zhang et al., 2024c), Wang et al. (2024a), CCoT (Cheng and Durme, 2024); etc. , leaf, text width=45.6em ] ] [  Scaling Law [  Brown et al. (2024), Snell et al. (2024), Wu et al. (2024c), Chen et al. (2024d); etc. , leaf, text width=45.6em ] ] [  Combination [  Marco-o1 (Zhao et al., 2024b), TTT (Akyürek et al., 2024), HiAR-ICL (Wu et al., 2024a); etc. , leaf, text width=45.6em ] ] ] ]
@@ 3.1 Updating the Model
In the inference stage, the ground-truth of test samples is unavailable. Thus many works attempt to design unsupervised or self-supervised objectives as learning signals. Existing learning signals can be classified into two categories based on whether the training process can be modified: test-time training (TTT) and fully test-time adaptation (FTTA). TTT assumes users can modify the training process by incorporating distribution-shift-aware auxiliary tasks. During test-time adaptation, the auxiliary task loss serves as the learning signal for optimization. Many self-supervised tasks have been shown to be effective as auxiliary tasks in image modality, such as rotation prediction (Sun et al., 2020), meta learning (Bartler et al., 2022), masked autoencoding (Gandelsman et al., 2022) and contrastive learning  (Liu et al., 2021; Chen et al., 2022). Among them, contrast learning has been successfully applied to test-time adaptation for visual-language tasks due to its generalization of self-supervised learning within and across modalities (Zhu et al., 2024).
In contrast, FTTA is free from accessing the training process and instead uses internal or external feedback on test samples as learning signals. Uncertainty is the most commonly learned signal, driven by the motivation that when test samples shift from the training distribution, the model’s confidence in its predictions is lower, resulting in higher uncertainty. Tent (Wang et al., 2021) uses the entropy of model predictions as a measure of uncertainty and updates the model by minimizing the entropy. MEMO (Zhang et al., 2022a) augments the data for a single test sample and then minimizes its marginal entropy, which is more stable compared to Tent in the single-sample TTA setting. However, minimizing entropy also has pitfalls, as blindly reducing prediction uncertainty may cause the model to collapse and make trivial predictions (Press et al., 2024; Zhao et al., 2023; Su et al., 2023a). Some works propose new regularization terms for minimizing entropy to avoid model collapse, including Kullback-Leibler divergence (Su et al., 2023a), moment matching (Hassan et al., 2023) and entropy matching (Bar et al., 2024). For specific tasks, a small amount of human feedback or external model rewards can also serve as high-quality learning signals. Gao et al. (2022) and Li et al. (2022b) utilize user feedback to adapt the QA model. Zhan et al. (2023) apply test-time adaptation to multilingual machine translation tasks by using COMET (Rei et al., 2020) for evaluating translation quality. In cross-modal tasks such as image-text retrieval and image captioning, RLCF (Zhao et al., 2024a) demonstrates its effectiveness by using CLIP scores as TTA signals.
@@
To advance the application of TTA in real-world scenarios, researchers must address challenges of efficiency and stability. To improve efficiency, many methods only fine-tune a small subset of parameters, such as normalization layers (Schneider et al., 2020; Su et al., 2023b), soft prompt (Shu et al., 2022; Hassan et al., 2023; MA et al., 2023; Feng et al., 2023; Niu et al., 2024), low-rank adaptation module (Imam et al., 2024), adapter module (Muhtar et al., 2024; Su et al., 2023a) and cross-modality projector (Zhao et al., 2024a). Although the number of parameters to fine-tune is reduced, TTA still requires an additional backward propagation. Typically, the time cost of a backward propagation is approximately twice that of a forward propagation. Thus, Niu et al. (2024) propose FOA, which is free from backward propagation by adapting soft prompt through covariance matrix adaptation evolution strategy.
The stability of TTA is primarily shown in two aspects. On the one hand, unsupervised or self-supervised learning signals inevitably introduce noise into the optimization process, resulting in TTA optimizing the model in the incorrect gradient direction. To address this, Niu et al. (2023) and Gong et al. (2024) propose noise data filtering strategies and the robust sharpness-aware optimizer. On the other hand, in real-world scenarios, the distribution of test samples may continually shift, but continual TTA optimization may lead to catastrophic forgetting of the model’s original knowledge. Episodic TTA (Wang et al., 2021; Shu et al., 2022; Zhao et al., 2024a) is a setting to avoid forgetting, which resets the model parameters to their original state after TTA on a single test sample. However, Episodic TTA frequently loads the original model, leading to higher inference latency and also limiting the model’s incremental learning capability. To overcome the dilemma, a common trick is the exponential moving average (Wortsman et al., 2022; Ye et al., 2022), which incorporates information from previous model states.

### v2

# Test-Time Compute: from System-1 Thinking to System-2 Thinking

@@ Abstract
The remarkable performance of the o1 model in complex reasoning demonstrates that test-time compute scaling can further unlock the model’s potential, enabling powerful System-2 thinking. However, there is still a lack of comprehensive surveys for test-time compute scaling. We trace the concept of test-time compute back to System-1 models. In System-1 models, test-time compute addresses distribution shifts and improves robustness and generalization through parameter updating, input modification, representation editing, and output calibration. In System-2 models, it enhances the model’s reasoning ability to solve complex problems through repeated sampling, self-correction, and tree search. We organize this survey according to the trend of System-1 to System-2 thinking, highlighting the key role of test-time compute in the transition from System-1 models to weak System-2 models, and then to strong System-2 models. We also point out a few possible future directions. [^1] https://github.com/Dereck0602/Awesome_Test_Time_LLMs.
Test-Time Compute: from System-1 Thinking to System-2 Thinking
Yixin Ji1, Juntao Li1thanks:   Corresponding author., Hai Ye2, Kaixin Wu3, Kai Yao3, Jia Xu3, Linjian Mo3, Min Zhang1 1School of Computer Science and Technology, Soochow University 2Department of Computer Science, National University of Singapore 3Ant Group jiyixin169@gmail.com; {ljt,minzhang}@suda.edu.cn yehai@comp.nus.edu.sg; {daniel.wkx,jiumo.yk,steve.xuj,linyi01}@antgroup.com
@@ 1 Introduction
Over the past decades, deep learning with its scaling effects has been the driving engine behind the artificial intelligence revolution. Particularly in the text modality, large language models (LLMs) represented by the GPT series (Radford et al., 2018, 2019; Brown et al., 2020; Ouyang et al., 2022; OpenAI, 2023) have demonstrated that larger models and more training data lead to better performance on downstream tasks. However, on the one hand, further scaling in the training phase becomes difficult due to the scarcity of data and computational resources (Villalobos et al., 2024); on the other hand, existing models still perform far below expectations in terms of robustness and handling complex tasks. These shortcomings are attributed to the model’s reliance on fast, intuitive System-1 thinking, rather than slow, deep System-2 thinking (Weston and Sukhbaatar, 2023). Recently, the o1 model (OpenAI, 2024), equipped with System-2 thinking, has gained attention for its outstanding performance in complex reasoning tasks. It demonstrates a test-time compute scaling effect: the greater the computational effort in the inference, the better the model’s performance.
@@
The concept of test-time compute emerged before the rise of LLMs and was initially applied to System-1 models (illustrated in Figure 1). These System-1 models can only perform limited perceptual tasks, relying on patterns learned during training for predictions. As a result, they are constrained by the assumption that training and testing are identically distributed and lack robustness and generalization to distribution shifts (Zhuang et al., 2020). Many works have explored test-time adaptation (TTA) to improve model robustness by updating parameters (Wang et al., 2021; Ye et al., 2023), modifying the input (Dong et al., 2024c), editing representations (Rimsky et al., 2024), and calibrating the output (Zhang et al., 2023c). With TTA, the System-1 model slows down its thinking process and has better generalization. However, TTA is an implicit slow thinking, unable to exhibit explicit, logical thinking process like humans, and struggles to handle complex reasoning tasks. Thus, TTA-enabled models perform weak System-2 thinking.
Figure 1: Illustration of test-time compute in the System-1 and System-2 model.
@@
Currently, advanced LLMs with chain-of-thought (CoT) prompting (Wei et al., 2022) have enabled language models to perform explicit System-2 thinking (Hagendorff et al., 2023). However, vanilla CoT is limited by error accumulation and linear thinking pattern (Stechly et al., 2024; Sprague et al., 2024), making it difficult to fully simulate non-linear human cognitive processes such as brainstorming, reflection, and backtracking. To achieve stronger System-2 models, researchers employ test-time compute strategies to extend model reasoning’s breadth, depth and accuracy, such as repeated sampling (Cobbe et al., 2021), self-correction (Shinn et al., 2023), and tree search (Yao et al., 2023). Repeated sampling simulates the diversity of human thinking, self-correction enables LLMs to reflect, and tree search enhances reasoning depth and backtracking.
To the best of our knowledge, this paper is the first to systematically review test-time compute methods and thoroughly explore their critical role in advancing models from System-1 to weak System-2, and ultimately to strong System-2 thinking. In Section 2, we present the background of System-1 and System-2 thinking. Section 3 and Section 4 detail the test-time compute methods for the System-1 and System-2 models. Then, we discuss future directions in Section 5. Additionally, we review benchmarks and open-source frameworks in Section 6.
@@ 2 Background
System-1 and System-2 thinking are psychological concepts (Kahneman, 2011). When recognizing familiar patterns or handling simple problems, humans often respond intuitively. This automatic, fast thinking is called System-1 thinking. In contrast, when dealing with complex problems like mathematical proofs or logical reasoning, deep and deliberate thought is required, referred as System-2 thinking—slow and reflective. In the field of artificial intelligence, researchers also use these terms to describe different types of models (LeCun, 2022). System-1 models respond directly based on internally encoded perceptual information and world knowledge without showing any intermediate decision-making process. In contrast, System-2 models explicitly generate reasoning processes and solve tasks incrementally. Before the rise of LLMs, System-1 models were the mainstream in AI. Although many deep learning models, such as ResNet, Transformer, and BERT, achieve excellent performance in various tasks in computer vision and natural language processing, these System-1 models, similar to human intuition, lack sufficient robustness and are prone to errors (Geirhos et al., 2020; Wang et al., 2022c; Du et al., 2023a). Nowadays, the strong generation and reasoning capabilities of LLMs make it possible to build System-2 models. Wei et al. (2022) propose the CoT, which allows LLMs to generate intermediate reasoning steps progressively during inference. Empirical and theoretical results show that this approach significantly outperforms methods that generate answers directly (Kojima et al., 2022; Zhou et al., 2023; Tang et al., 2024b; Feng et al., 2024a; Li et al., 2024h). However, current System-2 models represented by CoT prompting still have shortcomings. The intermediate processes generated by LLMs may contain errors, leading to cumulative mistakes and ultimately resulting in incorrect answers. Although retrieval-augmented generation (RAG) helps mitigate factual errors (Trivedi et al., 2023; Guan et al., 2024; Wang et al., 2024o; Ji et al., 2024), their impact on improving reasoning abilities remains limited. As a result, CoT-enabled LLMs are still at the weak system-2 thinking stage.
@@
 Auto-CoT (Zhang et al., 2023d), DAWN-ICL (Tang et al., 2024a); etc. , leaf, text width=34em ] ] ] [  Representation
@@
 Calibration [  $k$NN-MT (Khandelwal et al., 2021), AdaNPC Zhang et al. (2023c), Bi-$k$NN (You et al., 2024); etc. , leaf, text width=45.6em ] ] ] [  Test-time
@@
 GenRM (Zhang et al., 2024g), CriticRM (Yu et al., 2024b); etc. , leaf, text width=34em ] ] [ Verbal-based [  Liu et al. (2023), LLM-as-a-Judge (Zheng et al., 2023), Auto-J (Li et al., 2024b),
 Prometheus (Kim et al., 2024b, c), Fennec (Liang et al., 2024c); etc. , leaf, text width=34em ] ] ] [  Search
@@
 DiVeRSe (Li et al., 2023c), PRS (Ye and Ng, 2024)Zhang et al. (2024i); etc. , leaf, text width=34em ] [  Improvement training: ReST (Gulcehre et al., 2023), vBoN (Amini et al., 2024),
 BoNBoN(Gui et al., 2024), BOND (Sessa et al., 2024)Chow et al. (2024); etc. , leaf, text width=34em ] ] [ Self-correction [  Self-debug (Chen et al., 2024d), RIC (Kim et al., 2023), Critic (Gou et al., 2024),
 Shepherd (Wang et al., 2023b), MAD (Liang et al., 2024b), IoE (Li et al., 2024e),
 Refiner (Paul et al., 2024), Reflexion (Shinn et al., 2023), Du et al. (2023b); etc. , leaf, text width=34em ] [  Improvement training: GLoRe (Havrilla et al., 2024), SCoRe(Kumar et al., 2024),
 Self-correct (Welleck et al., 2023), Qu et al. (2024)Zhang et al. (2024k); etc. , leaf, text width=34em ] ] [ Tree Search [  ToT (Yao et al., 2023), RAP (Hao et al., 2023), rStar (Qi et al., 2024b),
@@
 MCTS-DPO (Xie et al., 2024), Zhao et al. (2024b)Zhang et al. (2024j); etc. , leaf, text width=34em ] ] ] ] [  Future
 Directions (§5) [  Generalization [  Jia (2024), GRM (Yang et al., 2024a), DogeRM (Lin et al., 2024b), Weak-to-strong (Burns et al., 2023); etc. , leaf, text width=45.6em ] ] [  Multi-modal [  MM-CoT (Zhang et al., 2024l), VoT (Wu et al., 2024b), Lee et al. (2024), LLaVA-CoT (Xu et al., 2024b); etc. , leaf, text width=45.6em ] ] [  Efficient [  Damani et al. (2024), OSCA (Zhang et al., 2024f), Wang et al. (2024e), CCoT (Cheng and Durme, 2024); etc. , leaf, text width=45.6em ] ] [  Scaling Law [  Brown et al. (2024), Snell et al. (2024), Wu et al. (2024c), Chen et al. (2024e); etc. , leaf, text width=45.6em ] ] [  Combination [  Marco-o1 (Zhao et al., 2024b), TTT (Akyürek et al., 2024), HiAR-ICL (Wu et al., 2024a); etc. , leaf, text width=45.6em ] ] ] ]
@@ 3.1 Updating the Model
In the inference stage, the ground-truth of test samples is unavailable. Thus, many works attempt to design unsupervised or self-supervised objectives as learning signals. Existing learning signals can be classified into two categories based on whether the training process can be modified: test-time training (TTT) and fully test-time adaptation (FTTA). TTT assumes users can modify the training process by incorporating distribution-shift-aware auxiliary tasks. During test-time adaptation, the auxiliary task loss serves as the learning signal for optimization. Many self-supervised tasks have been shown to be effective as auxiliary tasks in image modality, such as rotation prediction (Sun et al., 2020), meta learning (Bartler et al., 2022), masked autoencoding (Gandelsman et al., 2022) and contrastive learning  (Liu et al., 2021; Chen et al., 2022). Among them, contrast learning has been successfully applied to test-time adaptation for visual-language tasks due to its generalization of self-supervised learning within and across modalities (Zhu et al., 2024).
In contrast, FTTA is free from accessing the training process and instead uses internal or external feedback on test samples as learning signals. Uncertainty is the most commonly learned signal, driven by the motivation that when test samples shift from the training distribution, the model’s confidence in its predictions is lower, resulting in higher uncertainty. Tent (Wang et al., 2021) uses the entropy of model predictions as a measure of uncertainty and updates the model by minimizing the entropy. MEMO (Zhang et al., 2022a) augments the data for a single test sample and then minimizes its marginal entropy, which is more stable compared to Tent in the single-sample TTA setting. However, minimizing entropy also has pitfalls, as blindly reducing prediction uncertainty may cause the model to collapse and make trivial predictions (Press et al., 2024; Zhao et al., 2023; Su et al., 2023a). Some works propose new regularization terms for minimizing entropy to avoid model collapse, including Kullback-Leibler divergence (Su et al., 2023a), moment matching (Hassan et al., 2023) and entropy matching (Bar et al., 2024). For specific tasks, a small amount of human feedback or external model rewards can also serve as high-quality learning signals. Gao et al. (2022) and Li et al. (2022b) utilize user feedback to adapt the QA model. Zhan et al. (2023) apply test-time adaptation to multilingual machine translation tasks by using COMET (Rei et al., 2020) for evaluating translation quality. In cross-modal tasks such as image-text retrieval and image captioning, RLCF (Zhao et al., 2024a) demonstrates its effectiveness by using CLIP scores (Radford et al., 2021) as TTA signals. In language modeling, training with relevant contextual text at test time can reduce perplexity (Hardt and Sun, 2024; Wang et al., 2024l). Hübotter et al. (2025) theoretically shows that it reduces the uncertainty of test samples and proposes a more effective active learning selection strategy.
@@
To advance the application of TTA in real-world scenarios, researchers must address challenges of efficiency and stability. To improve efficiency, many methods only fine-tune a small subset of parameters, such as normalization layers (Schneider et al., 2020; Su et al., 2023b), soft prompt (Lester et al., 2021; Shu et al., 2022; Hassan et al., 2023; MA et al., 2023; Feng et al., 2023; Niu et al., 2024), low-rank module (Hu et al., 2022; Imam et al., 2024), adapter module (Houlsby et al., 2019; Muhtar et al., 2024; Su et al., 2023a) and cross-modality projector (Zhao et al., 2024a). Although the number of parameters to fine-tune is reduced, TTA still requires an additional backward propagation. Typically, the time cost of a backward propagation is approximately twice that of a forward propagation. Thus, Niu et al. (2024) propose FOA, which is free from backward propagation by adapting soft prompt through covariance matrix adaptation evolution strategy.
The stability of TTA is primarily shown in two aspects. On the one hand, unsupervised or self-supervised learning signals inevitably introduce noise into the optimization process, resulting in TTA optimizing the model in the incorrect gradient direction. To address this, Niu et al. (2023) and Gong et al. (2024b) propose noise data filtering strategies and the robust sharpness-aware optimizer. On the other hand, in real-world scenarios, the distribution of test samples may continually shift, but continual TTA optimization may lead to catastrophic forgetting of the model’s original knowledge. Episodic TTA (Wang et al., 2021; Shu et al., 2022; Zhao et al., 2024a) is a setting to avoid forgetting, which resets the model parameters to their original state after TTA on a single test sample. However, episodic TTA frequently loads the original model, leading to higher inference latency and also limiting the model’s incremental learning capability. To overcome the dilemma, a common trick is the exponential moving average (Wortsman et al., 2022; Ye et al., 2022), which incorporates information from previous model states.